a) Rationalise the denominator of
12/sqrt3
b) Work out the valule of
(sqrt2+sqrt8)^2
Please help :)

Answer by ewatrrr(24785)

12/√3 = 12√3 / 3 = 4√3
......
(√2 + √8) (√2 + √8) = 2 + 8 + 8 = 18
 
F First terms 2
O Outside terms √16 = 4
I Inside terms √16 = 4
L Last terms 8